The average depth of our rectangular swimming pool is 2 meters. If the pool
holds 520 m3
of water, and the length of one side is 6 meters less than twice
the length of the other side, what are the dimensions of our swimming pool?

Answer by ikleyn(52781)

.
The average depth of our rectangular swimming pool is 2 meters. If the pool
holds 520 m3 of water, and the length of one side is 6 meters less than twice
the length of the other side, what are the dimensions of our swimming pool?
~~~~~~~~~~~~~~~~~


Let w be the width, in meters;

then  (2w-6) meters is the length, according to the condition.


The equation for the volume is

    2*w*(2w-6) = 520.



Simplify ir step by step

    w*(2w-6)   = 520/2 = 260

    w*(w-3)            = 130.


At this point, the solution is OBVIOUS:  w = 13 meters.



You can get it algebraically, too

    w^2 - 3w - 130 = 0

    (w-13)*(w+10) = 0


The only positive root  w= 13 is the solution.


ANSWER.  The width is 13 meters;  the length is  2*13-6 = 26-6 = 20 meters.

Solved.